• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name E. Schulte

    Log entry time 15:58:05 on November 02, 2008

    Entry number 246267

    keyword=Day Shift Summary, November 2, 2008

    08:00 Shift starts with no beam, optics target in. MCC is working on an
    RF cavity.

    08:20 MCC called to let us know that they will be checking orbits,
    sending tune beam, and getting regular production beam back to the hall
    "soon".

    08:30 MCC is working on sending CW beam to the hall. We will check the
    beam positions with the BeO target (-2.0,0.3) on BPMA (-2.5,0.0) on BPMB.
    Check done at 2uA (our next production setting).

    09:06 MCC has been trying to send beam for about 30 minutes, but no luck.
    We are poised to check and restore the beam positions just as soon as
    they return production beam.

    09:13 MCC is going to try to send CW beam again.

    09:15 Start run on BeO Target with positons (-1.6, 0.15) BPMA; (-2.1,
    -0.1) BPMB. Spot++ was run. Run #: 2700 The beam spot on the BeO
    target looked pretty good.

    09:17 Switching to carbon hole target. Requested beam back with
    following beam positions (-1.5, 2.4) BPMA; (-1.5, 2.5) BPMB. Take run &
    check spot++. Run #: 2701

    09:43 OK we did a carbon hole with & without raster. Such is life.
    Raster on Run # 2702 & 2703

    10:01 Found that CODA had only 100K events set. Reset it to 1M. Request
    Harp Scan from MCC. Harps scanned: 1H03A/B.

    10:20 MCC reported that the results of the harp scan were 125um by 200um.
    The beam became unstable about the time we requested the harp scan.
    We'll get ready for the helium-3 target rate scans.

    10:25 Getting set up for the 3He target usage. The TO is performing the
    first NMR measurement.

    10:43 The beam has been really unstable for the last 30 minutes or so.
    The BigBite wire chambers, however, have been pretty quiet thus far.
    Even with all the beam tripping (before the NMR measurement got started).

    10:49 MCC called with a friendly reminder that we were not taking beam.
    I let them know that we were making the first NMR measurement, and would
    be getting back to them shortly.

    11:00 NMR Complete! Switching to optics target (BeO) for one last quick
    check (visual & spot++) of the beam position and then we'll move on to
    the 3He target for the raster rate scan.

    11:24 Spot++ check complete, time for the 3He cell.

    11:34 3He target in, requested 2uA from MCC with 3x3 raster.

    11:58 The raster scan results show no real difference between 3x3 & 4x4
    raster, so we'll go back to the 4x4 raster and proceed to the next NMR
    measurement. Then we'll take production data.

    12:00 Checking the NMR measurement. Changing target to NMR pickup coil
    position. The next NMR is needed at 14:00.

    12:04 Requested 5uA on 3He target with 4x4 raster for production running.
    Longitudinal Polarization Production. The BigBite wire chambers have
    been very well behaved so far.

    12:15 Started 5uA longitudinal production running. Raster 4x4 for 4 h of
    beam-on-target. To be taken in 30 minute increments.

    12:41 Believe it or not, we've had almost 20 whole minutes of relatively
    stable beam!

    13:13 Beam's been off for about 5-10 minutes at this point. Stopped the
    run. We'll start again when the beam comes back.

    13:22 Called MCC, they had a cavity drop and haven't been able to get it
    back up yet.

    13:31 HRS-L Q3 interlock fault. We will try to reset it.

    13:38 Called RC about interlock fault. We were unable to reset it
    remotely, so he asked me to call Tech On-Call (Scott) which I have done.

    13:44 Well, when it rains it pours. The ESR has a problem. This caused
    the liquid helium to fall below acceptable levels in the HRS-L Q3. I
    called MCC (that's how we found out about the ESR problem). They are
    going to call the cryo-coordinator. Meanwhile, back at the ranch, the TO
    took the opportunity to check the polarization via NMR. It's down to
    14%. He went to check the fibers. As luck would have it, we still have
    no beam.

    13:57 Well, the beam is back, but we still have our cryo and target
    problems. MCC is following up on the cryo problem.

    14:00 J.P. lets us know that we made a calculation error in the target
    polarization, so things are only half as bad as they seem.

    14:04 Updated RC on the status of Q3. He advised calling MCC every half
    hour until someone from Cryo Group gets in touch with us.

    14:08 At least the beam is no longer available...

    14:16 MCC called to let us know they had paged cryo to call us in the
    counting house. We still haven't heard from cryo.

    14:18 Cryo On-Call just called to let us know that he should be here in
    about 5 minutes.

    14:38 Called MCC. Cryo is looking at the problem but we don't have an
    estimate on how long it will take to fix it.

    14:46 Called Tech On-Call (Scot) to find out some info for Cryo On-Call.
    He determined that he needed to come in. Reported this to RC.
    Called for a Controlled Access with survey. Paged Alexandre.

    15:01 Just noticed that the liquid helium level on Q3 is now back to
    88.2%. I'm going to try to figure out what's going on.

    15:12 Q3 reset remotely and seems to be coming back online. We are
    cycling it and will reset it to the desired value.

    15:22 Back in business!

    15:48 Begin taking an NMR measurement. Beam has been crummy all day.